OpenStudy (anonymous):

Solve the system of equations by graphing x+y=-4 x+y=1

OpenStudy (anonymous):

Could you show me step by step?

OpenStudy (anonymous):

I don't know how to find the three points.

OpenStudy (displayerror):

The slope intercept form of a line is given by \[y = mx + b\] where m = slope and b = y-intercept. Put your two equations in slope intercept form: \[x + y = -4 \rightarrow y = -x - 4\] \[x + y = 1 \rightarrow y = -x + 1\] Two major points to look for are the y-intercept (where the line intersects the y-axis, ie. when x = 0) and the x-intercept (where the line intersects the x-axis, ie. y = 0) Using that idea, find the y-intercept and x-intercept and graph. Where the two lines intersects is your solution (at that point, the two lines have the same values for (x,y). Another way of solving would be to set your equations equal to each other. Notice above where I put the two equations in slope intercept form, both are equal to y. Since we're looking for the solution (where they intersect), at that point, the two lines must have the same x and y values. Using that fact, set them equal to each other: \[y = -x - 4 = -x + 1\]. Now you can determine your answer.

OpenStudy (anonymous):

I got y= 5 o.O

OpenStudy (displayerror):

The 'y' term isn't really there. I put it there because that's what both equations are equal to. The equation is actually -x - 4 = -x + 1. Solve this. As a check, plot them (by hand). Even when you get an answer, perform a sanity check: does it make sense that two straight lines of the same slope would intersect?

OpenStudy (anonymous):

x=5/2?

OpenStudy (anonymous):

2x-4=1 add 4 to both sides. 2x=5 divide by 2

OpenStudy (displayerror):

How do you get 2x? Solve this: \[-x-4=-x+1\]